What are the steps in solving this equation?
6-(n+3)=3n+5-2n

Answer by richwmiller(17219)

first calculate -(n+3) and remove the parentheses
-(n+3)=-n-3
so now we have
6-n-3=3n+5-2n
now consolidate the like terms
on the left we have two number s to be brought together.
6-3=3
and on the right we have 3n-3n
3n-2n=n
so now we have
3-n=n+5
now we want all numbers on one side and n on the other
so we add and subtract
we add n to both sides
3=2n+5
now we subtract 5 from both sides
-2=2n
now we divide by 2 (both sides)
-1=n
we are finished
we can check our work
6-(n+3)=3n+5-2n
plug -1 in for n
6-(-1+3)=3(-1)+5-2(-1)
let's get rid of the parentheses
6-2=-3+5+2
add and subtract
4=2+2=4 ok